OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g.

OpenStudy (anonymous):

F(x)=x^2 over (x^2 +4)

OpenStudy (anonymous):

@midhun.madhu1987 can you please help again :)

OpenStudy (perl):

g(x) = x^2 f(x) = x/(x+4) then f(g(x)) = x^2 / (x^2 + 4)

OpenStudy (anonymous):

Im sorry @perl i still dont understand

OpenStudy (anonymous):

how do express it as a composition of two functions?

OpenStudy (perl):

you can write f o g

OpenStudy (anonymous):

fog=x^2 over (x^2 +4) ?

OpenStudy (anonymous):

or something else?

OpenStudy (anonymous):

@ganeshie8 could you please help? :)

OpenStudy (anonymous):

@iGreen

OpenStudy (perl):

sorry, i was helping other people

OpenStudy (perl):

what was the question?

OpenStudy (anonymous):

i dont understand where you were going with the fog part. the question is express the following function F(x) as a composition of two functions f and g @perl

OpenStudy (anonymous):

F(x)=x^2 over (x^2 +4)

OpenStudy (perl):

we are breaking down F(x) into two functions, into a composition of two functions

OpenStudy (anonymous):

yea :)

OpenStudy (perl):

F(x) = f ( g ( x) ) , (caps matters here)

OpenStudy (anonymous):

ok

OpenStudy (perl):

so now, the 'inner' function, i chose to be x^2

OpenStudy (perl):

g(x) = x^2

OpenStudy (perl):

we can think of composition as two machines connected to each other

OpenStudy (anonymous):

alright, so g(x)=x squared

OpenStudy (anonymous):

for the first part

OpenStudy (perl):